rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#7602 — FS#11488 — DSLAM PPE75

Attached to Project— ADSL / SDSL
Incident
Backend / Core
CLOSED
100%
The DSLAM and the optic equipment of the site ppe75 no longer respond. We're investigating.
Date:  Friday, 26 September 2014, 15:47PM
Reason for closing:  Done
Comment by OVH - Friday, 29 August 2014, 13:10PM

Our technician is on site, they're investigating.


Comment by OVH - Friday, 29 August 2014, 13:22PM

We can't access the NRA, we'll try tomorrow morning.


Comment by OVH - Friday, 29 August 2014, 13:24PM

Our technician is on site.


Comment by OVH - Friday, 29 August 2014, 13:25PM

The electric problem on the optic equipment on the NRA is fixed. We will keep an eye on it.


Comment by OVH - Monday, 15 September 2014, 05:29AM

General electrical problem across the NRA pump PPE75.


Comment by OVH - Monday, 15 September 2014, 05:30AM

NRA is up, We will keep monitoring the equipment to confirm if it is working properly.